Q1b
Assume the aggregate supply relationship (between output and the price level) in a particular economy takes the following form, where all variables are represented as logs;
EQU-1' is; y2 = c - d (w - p)
where y2= level of output, p= price level, c & d are positive constants (i.e. c>0, d>0) w is the fixed value of the nominal wage and where c<dw.
Further assume that aggregate demand is determined as follows and subject to influence by fiscal & monetary policies as described by the following equation;
EQU-2' is; yd = e + a1,jg + a12 (m - p);
where yd is the level of aggregate demand, g represents the setting of government spending, m is the setting of nominal money supply, e is a positive (i.e. e>0), a1,j (where j=1,2) are coefficients indicating the input of each policy instrument on aggregate demand, and a1,j > 0 (where j=1,2)
The question is; using EQU-1 & EQU-2 above to solve for equilibrium values of p & y in terms of the settings of g and m, the a ('subscript' 1,j) and the other model
1
Expert's answer
2012-12-21T04:40:27-0500
EQU-1' is; y2 = c - d (w - p)
EQU-2' is; yd = e + a1,jg + a12 (m - p);
p = (e + a1,jg + a12m - c + dw) / (a12 - d)
where &quot;m&quot; quot;g&quot; are variable values,and &quot;p&quot; have a direct relationship, and other values are coefficients
and/or constants